OpenStudy (anonymous):

What is the slope of the line that passes through (10,0) and is parallel to y = (1/2)x + 3? ***Please explain!! :D

OpenStudy (anonymous):

Thanks for coming! :)

OpenStudy (anonymous):

IDK how to solve...

OpenStudy (anonymous):

can u pls explain and run me through how to solve so i understand? :D

OpenStudy (anonymous):

wouldnt you put the 10 in for x

OpenStudy (anonymous):

IDK!!! do u know why it would be like that??

OpenStudy (anonymous):

im trying to remember how to do these its been a while

OpenStudy (anonymous):

http://www.cpm.org/pdfs/state_supplements/Parallel_and_Perpendicular_lines.pdf this site might help

OpenStudy (anonymous):

0=(1/2)10+b b=-5 y=1/2x-5

OpenStudy (anonymous):

oh wait a sec!! Parallel lines have equal slopes right???? :O

OpenStudy (anonymous):

what i meant is.... you are correct when you said parallel lines have equal slopes....

OpenStudy (anonymous):

soooo the answer is that the slope is 1/2 am i correct???

OpenStudy (anonymous):